# Signal Processing over Time-Varying Graphs: A Systematic Review

### v1

@@ V-C Graph Lifelong Learning
VI Datasets, Evaluation Metrics and Applications
VI-A Datasets
Recently, with more and more attention have been paid to dynamic graph research, there are many time-varying datasets proposed to estimate the performance of different graph models. Here, we divide the time-varying graph datasets into three classes that include the Spatiotemporal Graphs (STGs) dataset, Discrete-Time Dynamic Graphs (DTDGs) dataset and Continuous-Time Dynamic Graphs (CTDGs) dataset. We summarize the statistics of these time-varying graph datasets in Table II.
1) Datasets for STGs In the spatiotemporal graphs research field, most of the representative datasets are used in the traffic domain, with one dataset use for neuroimaging. Los-loop [104] includes loop detector data collected from highways in Los Angeles. This dataset provides insights into traffic speed, volume, and flow on major LA highways, sampled typically at 5-minute intervals. PeMSD7 [15] dataset is part of the larger PeMS dataset from the California Department of Transportation (Caltrans) Performance Measurement System (PeMS), focusing on District 7 (Los Angeles) and includes speed, occupancy, and flow data from freeway sensors. It is used for traffic prediction and anomaly detection research. Traffic [113] consists of traffic data collected over a period of three months, with hourly granularity. It captures general traffic data across multiple urban and freeway road networks from 4,438 sensor stations located in the 7th District of California. METR-LA [106] contains data collected from sensors installed on highways in Los Angeles County, capturing traffic speed at various locations at 5-minute intervals. PeMS-Bay[106] is derived from the PeMS database and includes six months of traffic readings data from Bay Area highways. It provides traffic speed information at 5-minute intervals across numerous sensors. SZ-taxi [104] dataset contains GPS trajectories of taxis in Shenzhen, China, collected over several weeks. This dataset provides information on the location, speed, and timestamp of each taxi, offering a rich source of spatial-temporal traffic data in urban environments. Montevideo [130] captures one month of hourly traffic volume, speed, and congestion levels across various road segments of urban traffic data from the city of Montevideo, Uruguay. HCP [131] (Human Connectome Projectis) is a comprehensive collection of neuroimaging data designed to map the structural and functional connectivity of the human brain from over 1,200 healthy adult participants.
2) Datasets for CTDGs Since there are plenty of dynamic graph dataset for the continuous time dynamic graphs scenario, we choose some representative datasets in each domain to introduce them. Social Evolution [113] dataset comprises dynamic interactions between students at the Massachusetts Institute of Technology (MIT) Human Dynamics Lab. Collected through mobile devices, this dataset captures various forms of temporal social interactions, including phone calls, text messages, and physical proximity, over a period of months. Bitcoin-Alpha [134, 135] is a dynamic network of trust relationships between users on the Bitcoin Alpha platform, where users rate one another based on the trustworthiness of cryptocurrency transactions. LastFM [136] contains one month of who-listens-to-which song information, which was timestamped interactions between users and artists on the LastFM music streaming platform, capturing user preferences over time. UCI [138] is an online temporal communication, such as emails, between students from the University of California, Irvine, with edges in the graph corresponding to messages exchanged between users. Taobao [140, 141] collected from the Chinese e-commerce platform Taobao, contains temporal interactions between users and items, representing user behavior in an online marketplace, which links the meaning of the type of action (e.g., clicking, purchasing, or adding to a wishlist).
3) Datasets for DTDGs In the field of discrete-time dynamic graph research, the following datasets have become widely used, allowing researchers to analyze the evolution of relationships, dependencies, and interactions over discrete time intervals. UN Vote [143] captures voting records of countries in the United Nations General Assembly from 1946 onwards, with a node representing a country, and an edge between nodes reflecting similar voting behavior in a given year. Elliptic [147] contains Bitcoin transaction data from the Elliptic platform, with each transaction labeled as either ”licit”(e.g., exchanges, wallet providers, miners, licit services) or ”illicit”(e.g., scams, malware, terrorist organizations, ransomware, Ponzi schemes). Nodes represent Bitcoin addresses, and edges capture payment flow between these addresses in discrete time steps. HEP-TH [148, 149] is a citation network from arXiv’s High Energy Physics Theory section, where nodes represent papers and directed edges indicate citations between them. The dataset is segmented into monthly intervals, capturing the evolving structure of scientific collaboration and knowledge propagation over time. Twitter-Tennis [152] contains interactions on Twitter related to tennis events, where nodes represent Twitter users, and edges represent mentions, replies, or retweets within specific time intervals. Autonomous Systems [148] is a dynamic graph of Autonomous Systems (AS) over the internet, where nodes represent AS entities and edges indicate connections between them, captured at regular intervals. Flights [153] captures daily flight connections between airports, with nodes representing airports and directed edges indicating flights between them on a specific day. ABIDE [154, 155] (Autism Brain Imaging Data Exchange dataset) consists of functional MRI (fMRI) data from individuals diagnosed with autism spectrum disorder and neurotypical controls, which is often used to create brain connectivity networks over time, where nodes represent brain regions and edges represent functional connectivity.
VI-B Applications
On the application side, recent advancements in TV-GSP techniques promote the development of several fields, which include social science, biological analysis, transportation, environment, and finance due to the outstanding ability of dynamic graph representation in terms of capturing the dynamic relationship between irregularly distributed signals [156, 4, 5, 84].
1) Social science: In social science, time-varying graph signal processing (GSP) has emerged as a powerful tool for analyzing dynamic graph datasets, offering insights into evolving relationships, behaviors, and social interaction patterns. Applications in this domain span various areas, such as opinion dynamics, community detection, anomaly detection, and modeling social behavior.
Opinion Dynamics and Influence Propagation: TV-GSP is employed to study opinion dynamics and influence propagation across social networks, where individual nodes represent users or entities and edges represent their interactions over time. By treating these interactions as signals propagating across a dynamic network, TV-GSP enables the tracking of how opinions, behaviors, or sentiments spread and evolve. Enron [137] dataset serves as a prime example of how opinion and information diffusion can be tracked within an organization by analyzing dynamic interactions over time. This dataset captures email exchanges, offering a temporal sequence of internal communication and opinion exchanges among employees. Additionally, dynamic embedding techniques, such as Reddit and Wikipedia in [136] model user and entity behavior trajectories, allow predictions of future interaction and influence trends. By applying TV-GSP to these datasets, researchers can model the spread of information and identify patterns of influence within the corporate environment.
Community Detection: Community detection in dynamic networks benefits greatly from TV-GSP methodologies, which allow for the identification and tracking of evolving communities over time. In many social networks, communities are not static, and TV-GSP enables the monitoring of their changes, mergers, or splits. UCI in [138] provides an example by analyzing user engagement and behavioral patterns in an online setting. Similarly, [133] offers detailed, time-varying social interaction data that facilitates community evolution tracking and behavior analysis in physical and digital social contexts. This type of analysis, combined with TV-GSP, enables researchers to detect emerging communities and shifting user preferences over time, providing valuable insight into online social interactions.
Detection of Anomalous Patterns: TV-GSP is highly useful for anomaly detection in social networks, where unusual patterns in user behavior or interactions can be flagged in real-time. Such capabilities are crucial for monitoring social media for events like coordinated misinformation efforts or network breaches. In the financial sector, DGraphFin [139] dataset supports such TV-GSP applications, where anomaly detection algorithms are used to detect fraud and irregular transaction patterns. In social contexts, [152] introduces a centrality measure designed for continuously updating network streams, helping identify influential nodes and interactions as new data arrives. [157] employed a multi-level anomaly detection algorithm in dynamic graph settings, effectively detecting anomalies at different graph levels (e.g., node, subgraph) with interactive visualization. These methods prove valuable for anomaly detection within evolving social graphs where central actors can be continually monitored.
Modeling Social Behavior: Modeling social behavior during crisis events, such as natural disasters, pandemics, or public health emergencies, relies on TV-GSP to understand shifts in public sentiment, communication, and connectivity. [113] study exemplifies how temporal GSP techniques are used to assess the collective ”health state” or well-being of a community by tracking temporal patterns in communication and mobility data during significant events. This approach allows for deeper insights into behavior modeling, capturing the temporal evolution of individual behaviors and social influence mechanisms.
2) Biomedical analysis: Dynamic graph-based methods are particularly well-suited to representing temporally evolving interactions, such as neural connectivity patterns, gene expression, and physiological signals. Biomedical applications of TV-GSP encompass brain connectivity studies, predictive modeling in medical diagnostics, and protein interaction analysis.
Dynamic Functional Brain Connectivity Analysis: Functional connectivity analysis is an increasingly important brain function analysis modality, which regards connections as statistical codependency between the signals of different brain regions. TV-GSP has emerged as a valuable framework for analyzing functional brain connectivity, particularly in studies using time-series data from fMRI or EEG. Yu et al. [158] applied TV-GSP to assess dynamic brain connectivity in patients with schizophrenia by analyzing time-varying functional networks derived from fMRI data. This study highlighted how TV-GSP can identify connectivity alterations associated with mental disorders. Calhoun and Adali [159] further expanded by using TV-GSP techniques to track connectivity states across different brain regions, demonstrating applications in both healthy and clinical populations. Recently, Zhao et al. [155] by integrate particle filtering algorithm into graph neural network, achieves superior performance in brain disorder classification.
Medical Diagnostics Modeling: In the field of medical diagnostics, TV-GSP has enabled predictive modeling using dynamic patient data, offering tools for early diagnosis and outcome prediction. Xie et al. [160] introduced a graph convolutional neural network that predicts intensive care interventions based on temporal multivariate data from ICU records, demonstrating the potential of TV-GSP for real-time decision support in critical care environments. Another work by Li et al. [161] explored time-varying graph networks in EEG analysis to monitor cognitive health and diagnose neurological disorders, showing that dynamic graphs can reveal early-stage signs of brain dysfunction in clinical populations.
Protein Interaction and Structural Dynamics Analysis: TV-GSP provides a robust approach for studying dynamic protein interactions, enabling the modeling of temporal and structural changes within protein-protein interaction (PPI) networks. By applying TV-GSP, researchers can capture the evolution of interactions among proteins, which is essential for understanding functional pathways and responses to various stimuli. Song et al. [162] applied dynamic Bayesian networks to track protein signaling pathways over time, using time-varying graphs to model interactions during cellular processes. Furthermore, Li et al. [163] demonstrated how TV-GSP can track dynamic changes in PPI networks, highlighting structural adaptations in proteins under different environmental conditions. Recent work by Calazans et al. [164] leveraged TV-GSP for protein interaction analysis, using graph learning techniques to identify structural motifs and interaction dynamics in protein complexes.
3) Transportation: In transportation systems, TV-GSP has become an essential tool for analyzing dynamic datasets that capture the complex and evolving relationships between various components, such as vehicles, traffic signals, and road networks. Key applications include traffic flow prediction, congestion management, and route optimization.
Traffic Flow Prediction and Congestion Analysis: Traffic flow prediction is one of the most prominent applications of TV-GSP in transportation, where temporal GSP techniques analyze dynamic relationships between road segments or intersections to predict future traffic patterns. By treating traffic flow as a signal on a dynamic graph of road networks, GSP-based models capture both spatial dependencies (interconnectedness of adjacent road segments) and temporal variations (changes in flow over time). The widely used PeMS datasets from California, such as PeMSD7 [15] and PeMSBay [106], contain rich traffic sensor data, enabling effective applications of GSP for short- and long-term traffic prediction. In particular, [106] utilized GSP-based graph convolutional networks on the PeMS datasets to accurately forecast traffic flows by learning spatial-temporal dependencies across the network.
Transportation Network Analysis: TV-GSP also supports the analysis and optimization of public transportation networks, where interactions between transit nodes (e.g., bus stops, train stations) and connections (routes) change continuously over time. By analyzing passenger demand patterns, GSP methods can optimize transit scheduling and enhance network efficiency. For instance, [104] utilized GSP techniques on the SZ-Taxi dataset, which includes GPS-based movement data from taxis in Shenzhen, China, to analyze passenger demand patterns and recommend efficient resource allocation strategies. [130] further supports this application through the PyTorch Geometric Temporal library, which includes tools for implementing GSP methods that optimize public transit networks through spatial-temporal data analysis. In addition, [153] introduces a dataset from the OpenSky network, capturing real-time aircraft signals, which realizes monitoring the transportation condition in the context of air traffic and enables efficient airspace management.
Route Optimization and Navigation: For real-time route optimization, GSP models use dynamic graphs where edge weights vary with traffic conditions, road closures, or accidents, enabling route recommendations that adapt to evolving traffic patterns. Using datasets like METR-LA [106], which contains real-time traffic speed data from Los Angeles, [106] demonstrated that GSP-enabled models could dynamically adapt routing recommendations based on evolving traffic patterns, thus optimizing travel routes in real-time. Additionally, [113] provides a predictive framework for identifying probable path failures within evolving graphs, supporting route planning that can proactively avoid high-risk paths.
4) Environment: Since environmental data is often represented as dynamic graphs with evolving interactions among spatial or temporal nodes, TV-GSP can serve as a valuable framework for monitoring and modeling natural processes. The application of TV-GSP applied in environmental analysis including air quality forecasting, Meteorological prediction, and environmental hazard detection.
Air Quality Prediction: Due to the air quality is highly related to governments energy-saving and emission-reduction policies formulation and resident’s daily life outdoor activities and physical health, air quality prediction has became an critical research issue that attracts increasing attention. Air quality prediction requires the processing of spatially distributed, time-evolving data collected from networks of sensors across urban and rural landscapes. TV-GSP approaches enable the analysis and reconstruction of air pollution signals, providing real-time insights into pollutant levels. For example, [165] introduced a model for reconstructing time-varying graph signals applied to PM2.5 pollution data from California, enabling enhanced tracking of pollution levels over time by leveraging the spatial dependencies between monitoring stations. [166] developed a Sobolev smoothness-based framework that demonstrated effective signal recovery in environmental datasets, highlighting its application to dynamic air quality data. Another work [32] uses spatiotemporal smoothness, facilitating improved predictions of air quality by modeling correlations across sensor networks and adapting to temporal changes.
Meteorological Forecasting: TV-GSP also plays a crucial role in analyzing climate data, where dynamic graph models help capture seasonal and spatial variations. Ortega et al. [5] provided an overview of graph signal processing techniques tailored for environmental datasets, demonstrating their application to large-scale climate data with temporally and spatially varying features. Castro-Correa et al. [167] further advanced this area by developing Gegenbauer Graph Neural Networks specifically designed for reconstructing time-varying signals in complex climate data, allowing for the extraction of meaningful trends over extended periods. Moreover, Lou et al. [168] introduced high-order smoothness and adaptive low-rank techniques to handle the inherent complexity of climate signals, addressing challenges in handling missing data while maintaining the accuracy of climate trend predictions. There are also various meteorological prediction scenarios such as wind prediction [169], [170], [171] and temperature prediction [172], [173], illustrates the capability of TV-GSP in practical use.
Environmental Hazard Detection: Monitoring environmental hazards, such as floods, wildfires, and industrial pollution, benefits from the dynamic adaptability of TV-GSP frameworks, which allow for the real-time detection of anomalies and the tracking of evolving hazard patterns. Valdivia et al. [174] utilized wavelet-based TVGSP methods to visualize and process time-varying environmental data, enabling real-time analysis of environmental events such as wildfires. In a more recent study, Sun et al. [175] developed the Time-Varying Graph Convolutional Network (TVGCN) model to predict multivariate spatiotemporal series, particularly useful in forecasting hazard-related data such as rainfall and water levels in flood-prone areas.
5) Financial analysis: The integration of TV-GSP within financial analysis addresses diverse and complex challenges within dynamic financial systems, where interactions between entities such as stocks, transactions, and institutions change over time. TV-GSP approaches empower researchers and analysts to harness the temporal and structural dynamics of financial interactions, enabling real-time insights into evolving networks. Financial applications of TV-GSP include fraud detection, risk assessment, and stock price forecasting.
Fraud Detection and Anomaly Detection: TV-GSP enables anomaly detection in financial networks, which is critical for identifying irregular transaction patterns and preventing fraud. [135] provides a framework for detecting fraudulent users through the temporal analysis of user ratings, focusing on anomalies in user behaviors and their interactions within the network. Similarly, [147] leverages graph neural networks on a temporal dataset of Bitcoin transactions to detect and prevent money laundering activities in digital finance, leveraging the Bitcoin and Elliptic dataset to identify laundering activities across time-evolving Bitcoin networks.
Risk Assessment and Market Volatility Analysis: Predicting stock market volatility often requires monitoring dynamic, interdependent relationships across different stocks or sectors. The temporal dynamics captured through TV-GSP provide insights into correlations and dependencies that influence market risk and volatility. For instance, [176] employs a time-varying graph model to analyze the dynamic correlations in global stock indices, enabling enhanced forecasting capabilities for market shifts and volatility. In addition, [134] provides a method for predicting future edge weights in financial networks by analyzing both positive and negative relationships, such as cooperative and competitive interactions, within the market. These edge weight prediction techniques assist in understanding market volatility by focusing on how relationship strengths evolve, offering predictive insights into potential price fluctuations.
Stock Price Prediction: Stock price prediction is an essential component of financial trading, enabling traders to make informed decisions regarding the purchasing, selling, and retention of equities. Accurate predictions of future stock prices can help traders optimize their trading strategies and maximize their profits. By applying TV-GSP to financial networks, where nodes represent assets and edges represent correlations, these methods can capture and forecast changes in market conditions, which assists in stock price prediction and correlation analysis. Some common ways of building the TV-GSP for stock price prediction is to combine the graph neural networks (e.g., GCN, GAT) with time series predicting methods (e.g., RNN, LSTM) [177] [178]. While [146] examines the global agricultural trade network, by assessing how countries allocate cropland to domestic versus foreign demand, and considering multiple metrics, presenting a temporal graph of trade dependencies that evolve over time, thus enhancing decisionmaking surrounding trade.
VI-C Evaluation Metrics

### v2

@@ 6.3 Datasets for DTDG:
7 Conclusion and Future Direction
TVGs provide a powerful framework for modeling systems in which both signals and relational structures evolve over time. In this review, we examined the connections between TVGSP and TVGNNs, highlighting how theoretical signal-processing principles and data-driven learning models jointly contribute to the analysis of TVG data. By extending classical GSP tools to temporal graph settings and combining them with neural architectures capable of capturing nonlinear dependencies, these two paradigms provide complementary approaches for modeling complex real-world systems.
Despite rapid progress, research on TVGs is still in an early stage, and many theoretical and practical challenges remain. Continued integration between signal processing theory, machine learning, and emerging data-driven technologies is likely to open new directions for both methodology and applications. Building on the developments discussed in this review, several promising research directions are outlined below:
Theoretical Foundations and Scalability: The unified perspective between TVGSP and TVGNNs highlighted in this review provides a foundation for advancing the theoretical understanding of TVG models. Building on this connection, several fundamental questions remain important for the continued development of the field and pushing the frontiers of TVG research, particularly in terms of scalability, stability, convergence, and robustness of TVG-based algorithms. As graph sizes and time series lengths grow, scalable algorithms are needed to efficiently handle large-scale TVGs by reducing computational complexity without compromising performance [102]. Beyond STGs, TVGSP methods still lack mature theoretical formulations and efficient computational frameworks for DTDGs and especially CTDGs, partly due to the need to repeatedly handle topology evolution, irregular event dynamics, and potentially nonlinear dependencies. Developing frameworks that retain spectral interpretability while supporting dynamic and evolving structures remains a key challenge. Incorporating probabilistic or Bayesian frameworks may further enhance representation abilities and uncertainty quantification [32, 146, 98]. In parallel, a deeper understanding of phenomena such as over-smoothing and over-squashing in TVGNNs under dynamic graph conditions remains an important topic for future research.
Higher-Order Graph Signals: Beyond node-based representations, many real-world problems require modeling higher-order signals defined not only on nodes, but also on higher-order structures such as edges, faces, and multiway interactions, such as traffic flow on road segments [77], gene interaction dynamics [2], and multi-agent systems [61]. Higher-order structures, including hypergraphs, simplicial complexes, cell complexes, and multigraphs, provide a natural extension for such representations [11]. This has motivated the development of Topological Signal Processing, which generalizes GSP through the Hodge Laplacian to encode signals on simplicial or cell complexes beyond individual nodes [5, 145]. Using these operators, researchers have demonstrated edge-signal smoothing and denoising, embedding techniques for simplicial signals, semi-supervised edge learning, Hodge-wavelet analogues, localized topological Slepians, and FIR/convolutional filtering frameworks [6, 117, 162, 144, 8, 203, 159, 204, 82]. However, these techniques are still rarely explored in the context of STGs and remain largely unexplored for DTDGs and CTDGs. Collectively, these developments indicate that higher-order representations provide a principled pathway for modeling complex relational data, opening opportunities to extend current GSP and GNN approaches toward richer temporal, structural, and multi-entity signal domains [75].
Advancing Machine Learning Paradigms for TVGs: Learning from TVGs themselves remains a relatively new frontier in machine learning. Developing hybrid models that combine the interpretability of TVGSP with broader machine-learning paradigms can unlock new capabilities for learning from TVGs. While TVGNNs provide powerful neural architectures for dynamic graph learning, future research may also explore alternative learning frameworks such as probabilistic models, causal inference methods, and foundation-model–driven TVG analytics. Promising directions include multi-scale and multimodal learning strategies [67, 186], federated and privacy-preserving learning on TVGs, causal and counterfactual inference on dynamic networks, and machine unlearning for evolving graph systems [56, 57]. Beyond deterministic architectures, probabilistic and Bayesian learning frameworks offer promising directions for modeling uncertainty and capturing stochastic dynamics in TVGs. These developments may lead to more robust, generalizable, and explainable learning frameworks for dynamic graph analysis.
LLMs, AIGC, and Dynamic Graph Modeling: Recent advances in large language models (LLMs) and AI-generated content (AIGC) are beginning to intersect with TVG models, opening new opportunities for both TVGSP and TVGNN research [139]. On one hand, LLMs equipped with graph-augmented memory and reasoning modules provide a promising direction for learning high-level temporal and relational structure from dynamic graphs, potentially complementing classical spectral methods with semantic priors and interpretable temporal reasoning. On the other hand, the rapid rise of AIGC introduces new TVG application domains—such as evolving knowledge graphs, interaction networks in multi-agent generative systems, and continually updated recommendation environments—where graph topology and signals evolve at high frequency. These developments motivate hybrid TVG frameworks that integrate principled spectral modeling with data-driven temporal representation learning, and highlight emerging challenges related to scalability, continual adaptation, robustness, and reliability of TVGSP/TVGNN systems in LLM-driven environments.
Applications and Real-Time Adaptation: Exploring novel applications of TVGs in diverse and interdisciplinary domains can significantly propel the field forward. Beyond well-established areas, promising opportunities exist in neuroscience, epidemiology, smart cities, financial systems, cybersecurity, social dynamics, and bioinformatics. These applications benefit from the unique ability of over TVGs to represent and analyze highly dynamic, interconnected systems over time. Additionally, although current TVGSP has shown promising performance for real-time analysis on STGs, few TVGSP and TVGNN approaches extend this to the DTDGs and CTDGs. Future research should develop models capable of quickly adapting to evolving graph structures and signals in real time.
Overall, the integration of TVGSP principles with modern TVGNN frameworks highlights a promising paradigm for modeling and understanding complex time-varying systems. As theoretical foundations continue to mature and learning-based methods become more capable of capturing evolving relational structures, the synergy between signal-processing theory and data-driven machine learning is expected to play an increasingly important role in advancing both the methodology and practical applications of TVG analysis.
Acknowledgment
This work is supported by Shenzhen Science and Technology Innovation Commission Grant JCYJ20220530143002005, Shenzhen Ubiquitous Data Enabling Key Lab Grant ZDSYS20220527171406015, and Tsinghua Shenzhen International Graduate School Start-up Fund QD2022024C.
Appendix A Graph Acquisition and Representation Techniques
The formulation of an appropriate graph topology when TVG signals are present is crucial for the subsequent processing of both TVGSP and TVGNNs. Thus, it is necessary to examine key techniques for representing multivariate signals on topological structures.
The simplest and most common way of acquiring the graph topology is when the signals inherently exhibit or form a graph-like structure. In such cases, real-world data associated with the nodes can be constructed based on meaningful and realistic relationships. For example, naturally occurring graph structures appear in traffic forecasting tasks, where sensors placed along highways, recording data like traffic speed or flow, are modeled as nodes, and edges represent their spatial relationships [207]. In the case of meteorological data collection, variables such as longitude and latitude from several weather stations are aggregated and analyzed using a Gaussian Kernel to construct a k-nearest-neighbor graph [170]. A graph depicting the strength of 5G signal reception is created as detailed in [31] by documenting signal strength across various sensors. A slightly more abstract example of a real-world graph is a social network, where users are represented as nodes and their interactions form edges between them [59]. Another example is a citation network in which each node is a scientific paper and edges represent citations from one paper to another [163]. While many real-world datasets inherently exhibit graph structures, in other cases, the graph topology may not appear as straightforward. Specialized methods are employed to construct graphs that capture underlying relationships in the data.
Under the scenario where direct inference of the graph topology is not straightforward, a graph can still be formed. In statistical graphical models, statistical measures between two data elements are used to define the pairwise relationship. For example, the covariance matrix and the precision matrix can be used to form a graph because covariance is a statistical measure of the deviations between two variables, meaning that measuring and properly thresholding this pairwise relationship measure gives us a sparse graph [93]. The Gaussian Graphical Model is a specific example of a Statistical Graphical Model that represents a conditional dependency structure in a graph: the nodes of the graph are the multivariate Gaussian distributed variables, and edges between these variables are the conditional dependencies or non-zero partial correlation coefficients [210, 1]. GSP operations such as graph convolution, transforms, sampling, prediction, and regularization can be realized from a probabilistic and statistical aspect using the Gaussian Graphical Model [213]. The graph Laplacian can be learned using a merging of the Gaussian prior and a smoothness assumption on the multivariate signal, which is equivalent to learning the underlying topological structure [38]. A multivariate signal can be assumed to be a sparse linear combination of atoms from a dictionary based on a polynomial of graph Laplacians that represents localized patterns from a sparse prior [116]. Another sparse prior assumption topology learning approach can be found in [9]: the Gaussian Graphical Model is used in a way that data is trained in a neural network to learn a function calculating the empirical covariance matrices to form graph structures.
Graph embedding techniques leverage adjacency and topological properties to project signals and features onto an embedding space to facilitate the analysis, visualization, and application in downstream GSP tasks. In the spectral domain, graph embeddings can be achieved by leveraging the eigenvectors of graph representations such as the adjacency matrix [111]. In the spatial domain, early approaches to generating the graph embedding are to use the random walk to propagate through the graph and then embed the data on the walk trajectories as words using approaches such as word2vec [120]. The DeepWalk generates embeddings by performing random walks on the graph, treating each node sequence like a sentence in natural language processing [136]. The node2vec extended DeepWalk further by introducing a flexible search mechanism that balances breadth-first and depth-first search strategies [55]. Recently, a more common method to generate graph embeddings is through GNNs. The spectral GCN is achieved by feeding the spectral graph convolution in (4) into a nonlinear activation function and replacing the filter with learnable filters as if the filters are parameters learned from backpropagation [16]. The spatial GCN and the ChebNet are the spatial domain versions of GNNs with the spectral convolutions approximated by the Chebyshev polynomials, as shown in (7), and let the neural network learn the polynomial coefficient instead of the spectral filter [86, 34]. GraphSAGE exploits graph aggregations that sample and combine features from node neighborhoods, enabling an inductive framework where the model generalizes better to unseen nodes during inference [59]. The message-passing neural networks (MPNNs) further generalize the previous GNNs by introducing a flexible and localized method in which computations are defined at the node level based purely on the node local adjacencies instead of defining all computations at once [48]. These graph embedding techniques establish a foundation for integrating GSP with machine learning tasks, including classification and clustering, by transforming graph signals along with the information in the graph topology into representations suitable for non-graph machine learning models.
| Model | Wind Speed | Temperature | ||
| RMSE | MAE | RMSE | MAE | |
| GLMS | 2.43 ± 0.00 | 1.51 ± 0.00 | 1.30 ± 0.00 | 0.85 ± 0.00 |
| GNLMS | 2.97 ± 0.00 | 1.68 ± 0.00 | 1.21 ± 0.00 | 0.77 ± 0.00 |
| GNS | 3.03 ± 0.00 | 1.76 ± 0.00 | 1.69 ± 0.00 | 1.13 ± 0.00 |
| GCN | 2.52 ± 0.02 | 1.95 ± 0.01 | 2.76 ± 0.05 | 2.02 ± 0.01 |
| STGCN | 1.69 ± 0.41 | 1.27 ± 0.45 | 2.55 ± 0.12 | 1.92 ± 0.07 |
| RGDAN | 0.77 ± 0.03 | 0.50 ± 0.01 | 1.15 ± 0.10 | 0.79 ± 0.08 |
The pairwise relationships in graphs can be dynamic. Time-evolving processes over graphs can be captured by kernels formed by a preselected dictionary that fits the signal of interest, where partial correlation, partial Granger causality, non-linear structural equations, and vector autoregressions techniques are used along with assumptions of the graph signals, such as low-rank, sparsity, and smoothness [47, 154, 122]. A dynamic graph can be formed from the correlated properties, which are previously seen in static Statistical Graphical Models, in TVG signals by considering the space-time interactions in the underlying signal [108, 107, 195]. The topology inference on dynamic signals can be approached using a state-space formulation of a network process [28]. The Graphical LASSO is another instance of the Statistical Graphical Model method that estimates the inverse covariance matrix (precision matrix) with sparsity constraints enforced through an penalty. By combining Graphical LASSO with an ADMM-based algorithm, one can incorporate signals from neighboring time points and pairwise connections to learn the smoothly changing graphs [205]. Another smoothness-based graph topology learning task is to use the Sobolev smoothness as a measure to define a new smoothness function for TVG signals [49]. The work of Natali et al. proposes to use the structural equation model to learn graph topologies in an online fashion and further proposes a combination of the Gaussian graphical model, the structural equation model, and the smoothness-based model into an all-in-one online topology learning framework [127, 126]. The causal relationships between data can be modeled using a Vector Autoregressive model, making it possible to infer a graph topology from the underlying time-series vector [212]. A specific example of this approach can be seen in a time-varying gene activation graph formed by thresholding a regression matrix of a time-varying gene regression that reflects gene activation patterns over time, leading to a parametric representation of the time-evolution and influence between pairs of data [2]. Given an underlying graph structure at a one-time instance, assuming that the dynamics of a TVG only have some minor changes between time instances, a small perturbation analysis of the Laplacian matrix and its GFT can be done to model the time-evolution of the dynamic graph [160]. Topology inference of a dynamic topology from time-varying data can be tackled from another angle, using an iterative prediction-correction approach on a Gaussian Graphical Model [125].
Appendix B Evaluation Metrics